OpenStudy (anonymous):

Find the exact value of cos 2π/3.

OpenStudy (anonymous):

i have been stuck on this problem for like 10 min now

OpenStudy (anonymous):

pi/3 radians = 60 degrees, and cos is looking for the side adjacent to the 60 degree angle in a 30-60-90 triangle, the sides are in the ratio 1-sqrt(3)-2 so cos 60 = 1/2 2(pi/3) = 120 degrees-- the same triangle, except now, x is negative ==> cos 2pi/3 = -1/2

OpenStudy (anonymous):

dang i was so close i got -1/2 i was about to turn in my assignment

OpenStudy (anonymous):

gr8

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

you are aways wel come.